OpenStudy (mayankdevnani):

Simple question: A stone is dropped from building and 2 seconds later another stone is dropped. How far apart are these two stones by the time the first one has reached at a speed of 30 m/s. (Take g=10m/s^2) a) 80m b) 100m c) 60m d) 40m

OpenStudy (anonymous):

Well v=30m/s , u=0m/s , a=g=9.81m/s^2

OpenStudy (anonymous):

Then calculate t

OpenStudy (anonymous):

a=(v-u)/t

OpenStudy (mayankdevnani):

@AccessDenied @Algebraic! @Callisto @countonme123 @chandhuru @UnkleRhaukus @ajprincess @precal

OpenStudy (ghazi):

calculate the time taken by the first stone to reach 30 and then add 2 seconds in next calculation of your height for your second stone..and you are done

OpenStudy (mayankdevnani):

@sauravshakya a=10m/s^2

OpenStudy (ghazi):

a=g

OpenStudy (dls):

60m or 40m?

OpenStudy (mayankdevnani):

yaa 40 m is answer but how?

OpenStudy (dls):

i solved in mind :p @ghazi will explain

OpenStudy (mayankdevnani):

plz do it by step by step

OpenStudy (ghazi):

first calculate t @mayankdevnani @DLS please share the kudos :)

OpenStudy (dls):

you can use : v=u+at 30=0+10t t=3 seconds for first b

OpenStudy (mayankdevnani):

ok then

OpenStudy (dls):

then for 2nd ball it..says..t+2 seconds right? so t=5 seconds for 2nd ball.. find its speed

OpenStudy (dls):

you have both the speeds omg calculate distance o_O

OpenStudy (mayankdevnani):

ok then

OpenStudy (ghazi):

\[t=\frac{ V }{ a }=\frac{ 30 }{ 10 }\] then for the second stone \[h= 0.5 *g*t^2 \] but here t= 3+2 now calculate h for both the stones and find the difference

OpenStudy (dls):

that's what questionsays!!

OpenStudy (anonymous):

so the 2nd ball was dropped two seconds later than the first but falls for 5 seconds, whereas the first ball fell for 3 seconds. makes sense.

OpenStudy (mayankdevnani):

ok then

OpenStudy (mayankdevnani):

i want asnwer

OpenStudy (mayankdevnani):

@mathslover

OpenStudy (anonymous):

t=3 for first stone and t=3-2=1 for second stone

OpenStudy (anonymous):

yup

OpenStudy (dls):

A: H=0.5*10*9 B: H=0.5*10*1 45-5=40m :)

OpenStudy (ghazi):

use \[H= 0.5*g*t^2\] for both the timings... @DLS did it

OpenStudy (anonymous):